# SegIC: Unleashing the Emergent Correspondence for In-Context Segmentation

### v1

@@ 4.4 Ablation Study
In the ablation study, we report the performance on COCO-20 (in-domain one-shot segmentation), FSS-1000 (out-of-domain one-shot segmentation) and DAVIS-17 (zero-shot video object segmentation) to investigate the in-domain convergence and out-of-domain generalization capability.
Different vision foundation models. We further investigate the potential of each pre-trained VFM for in-context segmentation, including DINO-v1/v2 [6, 45], SAM-encoder [25], StableDiffusion-2.1 (SD-2.1) [52, 55], and CLIP [50]. For each VFM, we freeze it to extract image features and dense correspondences then map them into the latent space for mask decoding, equally. As shown in Tab. 2(a), the vision-language pre-trained model, CLIP, only achieves trivial performance on three datasets. This indicates that features learned by image-level vision-language alignment cannot be directly applied to fine-grained segmentation without fine-tuning. Additionally, its low-resolution pre-training may also contribute to poor performance. Meanwhile, other VLMs (i.e. DINO-v1/v2, SAM image encoder, and StableDiffusion) achieve decent performance on the segmentation benchmarks. Specifically, for StableDiffusion, we follow the usage in [52] to extract image features with 1-step diffusion. It is worth noting that the DINO-v2 series achieves the best performance, even outperforming the encoder of SAM, which is pre-trained for segmentation. We attribute this to its use of advanced vision contrastive learning as its pretext task, which learns the relation between different views within/across images, most relevant to dense correspondence. Besides, pre-training with high resolutions, e.g. DINO-v2, SD-2.1, may also help the segmentation. This implies that if the pretext task of a VFM is close to dense correspondence, this model is more likely to be easily extended into an in-context segmentation model.
Relation between correspondence and segmentation. We further conduct an experiment to study the relationship between correspondence and segmentation performance with different backbones. We evaluate a percent of correct keypoints score (PCK) on SPair-71k [41] following to [12], as well as the average segmentation score among COCO-20, FSS-1000, and DAVIS-17. As shown in Fig. 2, the performance on segmentation is proportional to correspondence, demonstrating a strong consistency between these two tasks. It further confirms our motivation that leverage the emergent correspondence for in-context segmentation.
@@
Ablations on in-context instructions. We study the importance of each component of in-context instructions (i.e. geometric, visual, and meta instructions). We conduct ablations on different combinations of components. As shown in Tab. 2(b), we find that the geometric and visual instructions tend to help the performance for out-of-domain segmentation and meta instructions (class name and task description in the ablation) benefit in-domain performance: when each component is used individually, the geometric and visual instructions obtain the best results on FSS-1000 and DAVIS-17, respectively. Meanwhile, meta instructions achieve the best performance on COCO-20, but with poor results on the other two datasets. Encouragingly, our method obtains the best performance among three tasks when using all three prompts, further demonstrating that our model can effectively transfer knowledge from in-context samples with the proposed in-context instructions.
@@
Ablations on training strategies. We investigate how the proposed training strategies affect performance. As shown in Tab. 2(c), since the two strategies, i.e. context reversion and negative entity augmentation, act as a form of “in-context augmentation”, they bring performance gains on COCO-20 and FSS-1000. However, the performance slightly drops on DAVIS-17. Furthermore, after combining them, our model achieves a consistent gain across all datasets, highlighting its efficacy.
@@
Dataset Combination. We further investigate the effectiveness of each dataset under our joint in-context training framework. We group the training data into two types: semantic segmentation (COCO, ADE20k) and instance segmentation (COCO, LVIS). As shown in Tab. 2(d), it is clear to see that when trained on COCO only, it achieves the best performance on COCO-20 with a weak performance on FSS-1000 and DAVIS-17. Based on this, we can enrich the training data from two directions: (1) use extra semantic segmentation data and (2) introduce instance segmentation tasks into training. For the former direction, after introducing ADE20k into training, the performance on FSS-1000 increases with little performance drop on the other two datasets. For the latter, we still use COCO as the training set but introduce instance segmentation with its instance-level annotation. It can be seen a clear performance boost of more than 20 points of score (45.170.1) on DAVIS-17, since video object segmentation needs an instance-level understanding. Finally, we further enrich our training data with LVIS [14], and our model achieves the most balanced performance among the three datasets.
@@ 5 Conclusion
We introduced SegIC, an end-to-end in-context segmentation framework that leverages the emergent correspondence of a single frozen vision foundation model. By training on standard segmentation datasets, SegIC achieved state-of-the-art performance on one-shot segmentation benchmarks, including COCO-20, FSS-1000, and LVIS-92. Impressively, SegIC demonstrated competitive performance on novel tasks, providing a cost-effective training approach for universal segmentation. In this work, our primary focus is on utilizing one in-context example per entity. In the future, we plan to explore utilizing multiple in-context examples to enhance contextual information. Additionally, we aim to investigate our model’s potential in instance-level segmentation, such as open-world instance segmentation. We do not anticipate any undesirable ethical or social impacts.
Acknowledgement We would like to express our appreciation for the valuable discussion with Xinlong Wang.

### v3

@@ 4.5 Ablation Study
In the ablation study, we report the performance on COCO-20i (in-domain one-shot segmentation), FSS-1000 (out-of-domain one-shot segmentation) and DAVIS-17 (zero-shot video object segmentation) to investigate the in-domain convergence and out-of-domain generalization capability.
geometric | visual | meta | COCO-20i | FSS-1000 | DAVIS-17 |
✓ | 68.8 | 87.3 | 65.8 | ||
✓ | 71.0 | 85.1 | 67.6 | ||
✓ | 73.7 | 63.7 | 21.6 | ||
48.9 | 58.1 | 21.1 | |||
✓ | ✓ | ✓ | 74.6 | 87.4 | 68.4 |
reversion | negative | COCO-20i | FSS-1000 | DAVIS-17 |
70.0 | 86.1 | 66.3 | ||
✓ | 72.9 | 86.9 | 65.0 | |
✓ | 73.1 | 86.5 | 64.3 | |
✓ | ✓ | 74.6 | 87.4 | 68.4 |
@@
Ablations on in-context instructions. We study the importance of each component of in-context instructions (i.e. geometric, visual, and meta instructions). We conduct ablations on different combinations of components. As shown in § 4.5, we find that the geometric and visual instructions tend to help the performance for out-of-domain segmentation and meta instructions (class name and task description in the ablation) benefit in-domain performance: when each component is used individually, the geometric and visual instructions obtain the best results on FSS-1000 and DAVIS-17, respectively. Meanwhile, meta instructions achieve the best performance on COCO-20i, but with poor results on the other two datasets. Encouragingly, our method obtains the best performance among three tasks when using all three prompts, further demonstrating that our model can effectively transfer knowledge from in-context samples with the proposed in-context instructions.
@@
Ablations on training strategies. We investigate how the proposed training strategies affect performance. As shown in § 4.5, since the two strategies, i.e. context reversion and negative entity augmentation, act as a form of “in-context augmentation”, they bring performance gains on COCO-20i and FSS-1000i with a slight drop on DAVIS-17. Furthermore, after combining them, our model achieves a consistent gain across all datasets, highlighting its efficacy.
semantic segmentation | instance segmentation | COCO-20i | FSS-1000 | DAVIS-17 | AVG | ||
COCOsem | ADE20k | COCOinst | LVIS | ||||
✓ | 76.3 | 80.9 | 45.1 | 67.4 | |||
✓ | ✓ | 75.6 | 82.3 | 40.1 | 66.0 | ||
✓ | ✓ | 75.7 | 83.5 | 70.1 | 76.4 | ||
✓ | ✓ | ✓ | ✓ | 74.6 | 87.4 | 68.4 | 76.8 |
@@
Dataset Combination. We further investigate the effectiveness of each dataset under our joint in-context training framework. We group the training data into two types: semantic segmentation (COCOsem, ADE20k) and instance segmentation (COCOinst, LVIS). As shown in Tab. 6, it is clear to see that when trained on COCOsem only, it achieves the best performance on COCO-20i but is weak on the other. Based on this, we can enrich the training data from two directions: (1) use extra semantic segmentation data and (2) introduce instance segmentation tasks into training. For the former direction, the performance on FSS-1000 increases after introducing ADE20k. For the latter, we still use COCO as the training set but introduce its instance-level annotation. It can be seen a clear performance boost on DAVIS-17 score (45.170.1), since video object segmentation requires instance-level understanding. Finally, we further enrich our training data with LVIS [16], achieving the best overall performance.
@@ 5 Conclusion
We introduced SegIC, an end-to-end in-context segmentation framework that leverages the emergent correspondence of a single frozen vision foundation model. By training on standard segmentation datasets, SegIC achieved state-of-the-art performance on one-shot segmentation benchmarks. Impressively, SegIC demonstrated competitive performance on novel tasks, providing a cost-effective training approach for universal segmentation. In this work, our primary focus is on utilizing one in-context example per entity. In the future, we plan to explore utilizing multiple in-context examples to enhance contextual information. Additionally, we aim to investigate our model’s potential in instance-level segmentation, such as open-world instance segmentation. We do not anticipate any undesirable ethical or social impacts.
Acknowledgement This project was supported by NSFC under Grant No. 62032006 and No. 62102092. We appreciate the valuable feedback from Xinlong Wang.
Appendix 0.A Additional Implementation Details
In this section, we provide additional details of the mask decoding and training pipeline.
Mask decoding. Inspired by recent query-based segmentation methods [7, 12, 11, 27], we utilize a lightweight query-based mask decoder to effectively map the in-context enhanced image features and object query to an output mask. We employ a learnable object query that will be used for the decoder’s output. The workflow of SegIC is concisely summarized in Pytorch-style pseudocode, as presented in Algorithm 1. Initially, the image feature and in-context instructions are projected into the same feature space for mask decoding. Subsequently, the projected in-context features are leveraged to enhance the image feature and object feature. The mask decoder then executes a multi-layered decoding process, structured as a four-step procedure within each layer, including (1) self-attention for the concatenated object query; (2) cross-attention from the image feature to the object query; (3) cross-attention from the query back to the image; and (4) calculating the mask. The mask calculation is performed using the image feature and the first element of the concatenated object feature, which corresponds to the position of the initial object query.
Training pipeline. In our main experiments, we adopt a mixed training scheme using both semantic segmentation datasets (COCO and ADE20k) and instance segmentation datasets (COCO and LVIS), as presented in Algorithm 2. We do not focus intensively on adjusting the dataset ratios, instead opting for uniform dataset-level sampling. For the segmentation datasets, we employ large-scale jittering (ranging from 0.1 to 2.0) for both the target image and in-context examples. These in-context examples are constructed based on the semantic class label of the target image, sampling one class per image during training. In the case of instance segmentation datasets, in-context examples are generated by applying two separate data augmentations to the same image. The instances from these differently augmented views then serve as mutual in-context examples. Our standard data augmentation techniques for this task include random resizing cropping (ranging from 0.3 to 1.0), random color jittering (with a 0.2 probability), and random horizontal flipping (with a 0.1 probability).
Appendix 0.B Additional Visualization
In this section, we provide more visualizations of the middle output and the predictions of SegIC.
Propagated mask. As outlined in Section 3.2, the propagated mask is derived from a weighted mean of dense correspondences according to the ground-truth mask of in-context samples. To facilitate visualization, we first apply the sigmoid function to map into . Subsequently, this range is transformed into RGB space using the JET color map. As depicted in Fig. 3, this process demonstrates that the propagated masks predominantly concentrate on the objects referenced in the in-context examples, providing strong guidance for the subsequent mask decoding process. This observation further demonstrates the emerging potential of pre-trained vision foundation models in the realm of segmentation tasks.
More qualitative results on VOS. We further provide more qualitative results on video object segmentation tasks (VOS). Note that SegIC is never trained on video datasets and just treats VOS as in-context segmentation using the first frame as in-context examples. As shown in Fig. 4, SegIC well handles challenging scenarios, including (a) occlusions, (b) interwoven objects, and (c) small objects.